# Signal-SGN: A Spiking Graph Convolutional Network for Skeleton Action Recognition via Learning Temporal-Frequency Dynamics

### v3

@@ III-B Multi-Scale Wavelet Transform Feature Fusion module
We first construct filters $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ based on Legendre polynomials $P_{k}(x)$ (details are provided in Appendix V-A) to decompose spiking-form features. The parameter $k$ specifies the number of frequency coefficients extracted, influencing the spectral resolution. We apply a linear (LN) transformation to the spiking features from the final layer $L$ of the backbone network to align the output tensor with the $k$ coefficients, ensuring compatibility with the filters.
$X_{\text{down}}=\operatorname{LN}\left(SN\left(X^{(L)}\right)\right)\in\mathbb{R}^{T\times k\times V}.$ (14)
To ensure compatibility with the wavelet transform, we verify whether the temporal dimension $T$ is a power of two. If not, $X_{\text{down}}$ is padded to the nearest power of two:
$X_{\text{extra}}=X_{\text{down}}[0:(2^{\lceil\log_{2}(T)\rceil}-T),:,:],$ (15)
$X_{\text{extended }}=\operatorname{concat}(X_{\text{down}},X_{\text{extra}},\text{ axis }=1),$ (16)
The extended spike-form feature $X_{\text{extended }}$ is decomposed by the filter matrices $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ into detail coefficients $D_{j}$, capturing high-frequency detailed features, and scaling coefficients $S_{j}$, retaining low-frequency trend features. The Algorithm 2 in Appendix V-B2 details the complete process.
@@
To enhance the interaction between detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ across decomposition levels $j$, we redefine the spiking cross-attention operation as:
$D_{j}^{\text{out}}=\left(\Phi_{Q}(D_{j})\cdot\Phi_{K}(D_{j})^{T}\right)\Phi_{V}(S_{j})*s,$ (17)
$S_{j}^{\text{out}}=\left(\Phi_{Q}(S_{j})\cdot\Phi_{K}(S_{j})^{T}\right)\Phi_{V}(D_{j})*s,$ (18)
where the temporal resolution is halved at each level. At the final level $J$, the scaling coefficient $S_{J}$ bypasses cross-attention and is directly upsampled to match the temporal resolution $T/2$:
$\hat{S}_{J}=\text{Upsample}(S_{J},T/2).$ (19)
All previous levels’ outputs $\{D_{j}^{\text{out}},S_{j}^{\text{out}}\}_{j=0}^{J-1}$ are similarly upsampled, and the aggregated feature map is computed as:
$X_{\text{MLO}}=\operatorname{GAP}\left(\Phi\left(\sum_{j=0}^{J-1}\hat{D}_{j}+\sum_{j=0}^{J}\hat{S}_{j}\right)\right).$ (20)
@@
The final feature map $X_{\text{MLO}}$ integrates dynamic spiking features and spectral characteristics, which are fused with the output $X^{(L)}$ from the last backbone layer undergoing $SN$ and GAP. Then the fused feature proceed to a fully connected (FC) layer for classification:
$\hat{y}=\operatorname{FC}\left(\operatorname{GAP}\left(SN\left(X^{(L)}\right)\right)+\beta X_{\text{MLO}}\right),$ (21)
where $\beta$ is a learnable parameter. Finally, the cross-entropy loss is employed to optimize the classification task.
IV Experiments
We evaluate the Signal-SGN model on three skeleton-based action recognition datasets: NTU-RGB+D [30], NTU-RGB+D 120 [31], and NW-UCLA [32]. Detailed descriptions of the datasets, along with hyperparameter design, FLOPs, synaptic operations (SOPs), and theoretical energy consumption calculations, are provided in the Appendix V-C, V-D and V-E.
IV-A Comparison with the State-of-the-Art
TABLE I: Comparative results on NTU RGB+D and NTU-RGB+D 120. We evaluate our model in terms of classification accuracy (%). Xs and Xv represent cross-subject and cross-view splits.
Model (ANN/SNN) Param. (M) FLOPs (G) SOPs (G) Power (mJ) Xs (60) Xv (60) Xs (120) Xv (120)
Part-aware LSTM [33] - - - - 62.9 70.3 25.5 26.3
Spatio-Temporal LSTM [34] - - - - 61.7 75.50 55.7 57.9
ST-GCN [6] 3.10 3.48 - 16.01 81.5 88.3 70.7 73.2
2S-AGCN [7] 3.48 37.32 - 182.87 88.5 95.1 82.5 84.3
Shift-GCN [5] - 10.00 - 46.0 90.7 96.5 85.3 86.6
MS-G3D [35] 3.19 48.88 - 239.51 91.5 96.2 86.9 88.4
CTR-GCN [9] 1.46 7.88 - 36.25 92.4 96.8 88.9 90.6
Spikformer [13] ICLR 2023 4.78 24.07 1.69 2.17 73.9 80.1 61.7 63.7
Spike-driven Transformer [14] NIPS 2023 4.77 23.50 1.57 1.93 73.4 80.6 62.3 64.1
Spike-driven Transformer V2 [15] ICLR 2024 11.47 38.28 2.59 2.91 77.4 83.6 64.3 65.9
Spiking Wavelet Transformer [29] ECCV 2025 3.24 19.3 1.48 2.01 74.7 81.2 63.5 64.7
MK-SGN [17] 2.17 7.84 0.68 0.614 78.5 85.6 67.8 69.5
Signal-SGN (Joint) 1.74 1.62 0.314 0.372 80.5 87.7 69.2 72.1
Signal-SGN (Bone+Joint) 1.74 3.24 0.628 0.744 82.5 89.2 71.3 74.2
Signal-SGN (4 ensemble) 1.74 6.48 1.28 1.488 86.1 93.1 75.3 77.9
The Param. represents the model size for the network itself. For ANN, the metrics such as FLOPs, power, and accuracy are reported and calculated based on the maximum ensemble of modalities as specified in the referenced paper.
TABLE II: Comparative results on NW-UCLA. We evaluate our model in terms of classification accuracy (%).
Model Param. FLOPs SOPs Power Acc
(ANN/SNN) (M) (G) (G) (mJ) (%)
Ensemble TS-LSTM [3] - - - - 89.2
DC-GCN+ADG [18] 2.48 3.6 - 16.56 95.3
CTR-GCN [9] 1.46 2.32 - 10.67 96.5
Spikformer [13]ICLR 2023 4.78 8.55 0.513 0.673 85.4
Spike-driven Transformer [14]NIPS 2023 4.77 8.22 0.501 0.643 83.4
Spike-driven Transformer V2 [15]ICLR 2024 11.47 11.96 0.717 0.841 89.4
Spiking Wavelet Transformer [29]NIPS 2023 3.24 7.36 0.438 0.573 86.7
MK-SGN [17] 2.17 3.66 0.165 0.207 92.3
Signal-SGN (Joint) 1.74 0.67 0.161 0.147 92.7
Signal-SGN (Bone+Joint) 1.74 1.24 0.322 0.294 93.1
Signal-SGN (4 ensemble) 1.74 2.48 0.644 0.588 95.9
Several SOTA methods have adopted multi-stream fusion strategies [7, 5, 9]. For a fair comparison, we implement the same four-stream fusion approach for our Signal-SGN, incorporating ”joint”, ”bone,” ”joint motion”, and ”bone motion.” We compare Signal-SGN with state-of-the-art methods across three datasets. To compare with SOTA SNN models[13, 14, 15, 29], we perform unimodal reproductions using the same hierarchical architecture as MK-SGN[17], demonstrating the superiority of our method. The parameters of our Signal-SGN model, including FLOPs, SOPs, energy consumption, and accuracy for both single-modal and multi-modal scenarios, are presented in Tables I and II. In the NTU-RGB+D dataset, we resize the skeleton sequences to 16 frames, our model outperforms SOTA SNN models in accuracy while significantly reducing FLOPs, SOPs, and theoretical energy consumption when using single-modality joint data. Furthermore, compared to classical ANN models, our model achieves a higher accuracy than ST-GCN when integrating four modalities, which is quite rare for SNN. Our model also requires fewer FLOPs than the lightweight CTR-GCN and Shift-GCN models. In terms of theoretical energy consumption, our model reduces the energy usage to less than 5% compared to GCN-based methods, which is highly significant in practical applications.
In the NW-UCLA dataset, we also resize the skeleton sequences to 16 frames. Our model achieves an accuracy of 95.9%, which is only 0.6% lower than the classic CTR-GCN model while reducing theoretical energy consumption to 5.51% of CTR-GCN’s. This further demonstrates the effectiveness and energy-saving capabilities of our model. Additionally, removing the extra spike time step dimension reduces FLOPs and improves memory efficiency, making the model more trainable on standard devices compared to the previous SGN model.
IV-B Ablation Studies
@@
To evaluate the contributions of the primary components in Signal-SGN, we conducted ablation studies on the NTU-RGB+D (cross-subject) dataset using joint data, with all skeleton sequences resized to 16 frames. Table III summarizes classification performance for different main module configurations. Appendix V-F presents a detailed analysis.
@@
TABLE III: Comparison of experimental results under different module settings
Method 1D-SGC FSC MWTF Param. FLOPs/SOPs Acc
(M) (G) (%)
Ours - - 1.03 1.12/0.252 70.5
- 1.65 1.60/0.312 78.3
1.74 1.62/0.314 80.5
When only the 1D-SGC module is used, the model achieves an accuracy of 70.5%, demonstrating its ability to extract spiking-form graph features while maintaining low parameter count (1.03M) and computational costs (1.12G FLOPs/0.252G SOPs). Introducing FSC module significantly improves accuracy to 78.3%, highlighting the role of frequency-domain feature extraction in enhancing representations. The addition results in a moderate increase in parameter count to 1.65M and computational costs to 1.60G FLOPs/0.312G SOPs. Finally, incorporating the MWTF module further boosts accuracy to 80.5%. The improvement comes with a marginal parameter increase (1.74M) and computational costs (1.62G FLOPs/0.314G SOPs), demonstrating the MWTF module’s effectiveness in capturing multi-scale frequency features and fusing temporal and spectral information. These results validate the complementary roles of the 1D-SGC, FSC, and MWTF modules. The 1D-SGC module focuses on temporal-domain graph features, while the FSC module captures frequency-domain information. The MWTF module further enhances the model by integrating multi-scale frequency features, establishing Signal-SGN as an efficient framework for skeleton-based action recognition.
V Conclusion

### v5

@@ 3.4. Multi-Scale Wavelet Transform Feature Fusion module
We first construct filters $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$ based on Legendre polynomials $P_{k}(x)$ (details are provided in Appendix B) to decompose spiking-form features. The parameter $k$ specifies the number of frequency coefficients extracted, influencing the spectral resolution. We apply a linear transformation to the spiking features from the final layer $L$ of the backbone network to align the output tensor with the $k$ coefficients, ensuring compatibility with the filters.
(13) $X_{\text{down}}=\operatorname{LN}\left(SN\left(X^{(L)}\right)\right)\in\mathbb{R}^{T\times k\times V}.$
To ensure compatibility with the wavelet transform, we first check whether the temporal dimension $T$ is a power of two. If it is not, we extend the input spike-form feature $X_{\text{down}}$ by concatenating extra frames, denoted as $X_{\text{extra}}$, which are duplicated copies of the beginning frames. The total number of time steps is then increased to the nearest power of two. Formally, the number of extra frames added is calculated as $2^{\lceil\log_{2}T\rceil}-T$, and the final extended input becomes $\widetilde{X}=[X_{\text{down}},X_{\text{extra}}]$ along the temporal axis.
The extended input $\widetilde{X}$ is then recursively decomposed by a set of wavelet filter matrices, namely $\Lambda_{0}$, $\Lambda_{1}$, $\Gamma_{0}$, and $\Gamma_{1}$, yielding two types of outputs at each level: detail coefficients $D_{j}$, which capture high-frequency, localized temporal features, and scaling coefficients $S_{j}$, which preserve the lower-frequency trend components, as shown in Figure 4. A complete iterative decomposition process is provided in Algorithm 2 in Appendix C.2.
Figure 4. The process of DWT
@@
To enhance the interaction between detail coefficients $D_{j}$ and scaling coefficients $S_{j}$ across decomposition levels $j$, we first upsample them to the original temporal resolution $T$. Then, a spiking cross-attention mechanism is applied to capture the mutual dependencies between high-frequency and low-frequency components:
(14) $\hat{D}_{j}=\left(\Phi_{Q}(D_{j})\cdot\Phi_{K}(D_{j})^{T}\right)\Phi_{V}(S_{j})\cdot s,$
(15) $\hat{S}_{j}=\left(\Phi_{Q}(S_{j})\cdot\Phi_{K}(S_{j})^{T}\right)\Phi_{V}(D_{j})\cdot s,$
where $s$ is a learnable scaling factor. The scaling coefficient from the final level $S_{J}$ is only upsampled (without interaction) and denoted as $\hat{S}_{J}$. All outputs $\{\hat{D}_{j},\hat{S}_{j}\}_{j=0}^{J-1}\cup\{\hat{S}_{J}\}$ are aggregated and passed through spiking activation and global average pooling to yield the multi-level output:
(16) $\hat{X}=\operatorname{GAP}\left(\Phi\left(\sum_{j=0}^{J-1}\hat{D}_{j}+\sum_{j=0}^{J}\hat{S}_{j}\right)\right).$
@@
The resulting feature $\hat{X}$ is then fused with the SN-processed output of the final backbone layer $X^{(L)}$, and the fused representation is sent to a fully connected layer for classification:
(17) $\hat{y}=\operatorname{FC}\left(\operatorname{GAP}\left(SN\left(X^{(L)}\right)\right)+\beta\hat{X}\right),$
where $\beta$ is a learnable parameter balancing the contribution of temporal and spectral features. Cross-entropy loss is used for optimization during training.
4. Experiments
To demonstrate the effectiveness of Signal-SGN, we conduct skeleton-based action recognition experiments on three large-scale datasets. We compare our model against strong ANN and SNN baselines, and perform ablation studies to evaluate the contribution of each module. The following sections cover the experimental setup in detail, including dataset settings, hyperparameter configurations, comparisons with state-of-the-art methods, and ablation studies.
4.1. Dataset
Table 1. Comparison of Signal-SGN with prior ANN- and SNN-based models on NTU RGB+D, NTU RGB+D 120, and NW-UCLA datasets. We report classification accuracy (%), model size (Param.), computational cost (FLOPs and SOPs), and energy consumption (Power).
Model (ANN/SNN) Param. FLOPs SOPs Power NTU RGB+D NTU RGB+D 120 NW-UCLA
(M) (G) (G) (mJ) Xs(%) Xv(%) Xs(%) Xt(%) (%)
Part-aware LSTM (Du Y., 2015) - - - - 62.9 70.3 25.5 26.3 -
Spatio-Temporal LSTM (Liu et al., 2016) - - - - 61.7 75.5 55.7 57.9 -
ST-GCN (Yan et al., 2018) 3.10 3.48 - 16.01 81.5 88.3 70.7 73.2 -
2S-AGCN (Shi et al., 2019) 3.48 37.32 - 182.87 88.5 95.1 82.5 84.3 -
Shift-GCN (Cheng et al., 2020b) - 10.00 - 46.00 90.7 96.5 85.3 86.6 94.6
MS-G3D (Liu et al., 2020) 3.19 48.88 - 239.51 91.5 96.2 86.9 88.4 -
CTR-GCN (Chen et al., 2021) 1.46 7.88 - 36.25 92.4 96.8 88.9 90.6 96.5
Spikformer (Zhou et al., 2022) 4.78 24.07 1.69 2.17 73.9 80.1 61.7 63.7 85.4
Spike-driven Transformer (Yao et al., 2024b) 4.77 23.50 1.57 1.93 73.4 80.6 62.3 64.1 83.4
Spike-driven Transformer V2 (Yao et al., 2024a) 11.47 38.28 2.59 2.91 77.4 83.6 64.3 65.9 89.4
Spiking Wavelet Transformer (Fang et al., 2025) 3.24 19.30 1.48 2.01 74.7 81.2 63.5 64.7 86.7
MK-SGN (Zheng et al., 2024) 2.17 7.84 0.68 0.614 78.5 85.6 67.8 69.5 92.3
Signal-SGN (Joint) 1.74 1.62 0.314 0.372 80.5 87.7 69.2 72.1 92.7
Signal-SGN (Bone+Joint) 1.74 3.24 0.628 0.744 82.5 89.2 71.3 74.2 93.1
Signal-SGN (4 ensemble) 1.74 6.48 1.28 1.488 86.1 93.1 75.3 77.9 95.9
Param. represents the model size for the network itself. For ANN methods, FLOPs, power, and accuracy are reported under standard protocols. For SNN methods, SOPs and energy estimates are computed based on spike-driven inference.
NTU RGB+D (Shahroudy et al., 2016) is a large-scale benchmark dataset for 3D human action recognition, consisting of 56,880 skeleton-based video samples from 60 action classes. Each sample contains 3D coordinates of 25 human body joints captured by the Kinect v2 sensor. We follow the standard evaluation protocols, including cross-subject (Xs)—where training and testing subjects do not overlap—and cross-view (Xv)—where training and testing samples are captured from different camera angles. NTU-RGB+D 120 (Liu et al., 2019) extends NTU 60 to 120 action categories and 114,480 video samples, with more subjects and greater view diversity. Evaluation is performed under cross-subject (Xs) and cross-setup (Xt) protocols, where the latter refers to training and testing samples being collected under different environmental and camera configurations, posing a greater challenge for model generalization. NW-UCLA (Wang et al., 2014) contains 10 action classes and 1,494 samples captured from three different viewpoints. Each sample includes 3D skeletons with 20 joints. The dataset is designed to test model robustness under cross-view scenarios.
4.2. Model Implementation and Configuration
Our model is implemented using PyTorch [^1] https://github.com/rwightman/pytorch-image-models and SpikingJelly [^2] https://github.com/fangwei123456/spikingjelly, and all experiments are conducted on four NVIDIA V100 GPUs. The backbone network consists of 4 layers with input and output channel sizes of 3 and 64, 64 and 64, 64 and 128, 128 and 256. We use SGD optimizer with a learning rate of 0.1, decay rate of 0.1 at step 110, batch size of 64, training for 150 epochs with a dropout rate of 0.3, weight decay of 0.0001, and auxiliary loss weight $\omega$ set to 0.03.
4.3. Comparison with the State-of-the-Art
Several SOTA methods adopt multi-stream fusion strategies for improved performance (Shi et al., 2019; Cheng et al., 2020b; Chen et al., 2021). To ensure a fair comparison, we employ the same four-stream fusion protocol in Signal-SGN, incorporating joint, bone, joint motion, and bone motion modalities. We evaluate Signal-SGN against both ANN and SNN baselines on NTU RGB+D, NTU RGB+D 120, and NW-UCLA. Classification accuracy, model size, FLOPs, synaptic operations(SOPs), and estimated energy consumption are summarized in Table 1, with calculation details provided in Appendix A. On the NTU RGB+D and NTU RGB+D 120 datasets, all sequences are downsampled to 16 frames. Compared to prior SNNs, Signal-SGN (Joint) achieves 80.5% accuracy on NTU RGB+D (Xs) while requiring only 0.372 mJ of energy—an 82.9% reduction compared to Spikformer’s 2.17 mJ, and a 98.9% reduction compared to the ANN-based MS-G3D (239.51 mJ). Notably, our four-stream fused model further boosts accuracy to 86.1% (Xs) and 93.1% (Xv) on NTU RGB+D and 75.3% (Xs) / 77.9% (Xt) on NTU RGB+D 120, while still keeping energy usage at 1.488 mJ, which is over 96% lower than CTR-GCN (36.25 mJ) and 99.4% lower than MS-G3D. This demonstrates Signal-SGN’s superior trade-off between performance and efficiency. On the NW-UCLA, Signal-SGN achieves a competitive 95.9% accuracy, only 0.6% below CTR-GCN, while using less than 5% of its theoretical energy—further validating the model’s generalizability and low-cost design.
4.4. Ablation Studies
@@
To evaluate the contributions of the primary components in Signal-SGN, we conducted ablation studies on the NTU-RGB+D (Xs) dataset based on joint data, with all sequences resized to 16 frames.
Table 2. Module-wise ablation study on NTU-RGB+D (Xs). We report model size, computational cost, and top-1 accuracy. Spikformer is included as a baseline for external comparison.
3D C.N. 1D-SGC FSC MWTF Param. (M) FLOPs / SOPs (G) Acc (%)
Spikformer (baseline)(Zhou et al., 2022) 4.78 24.07 / 1.69 73.9
- - - 1.03 1.12 / 0.252 69.7
- - 1.65 1.60 / 0.312 77.0+7.3
- 1.74 1.62 / 0.314 79.3 +2.3
- - 1.03 1.12 / 0.252 70.5
- 1.65 1.60 / 0.312 78.3+7.8
1.74 1.62 / 0.314 80.5+2.2
4.4.1. Module-wise Ablation Analysis
We first module-wise ablation analysis to evaluate the effectiveness of the proposed components in Signal-SGN. As summarized in Table 2, the base model with only the 1D-SGC module achieves 69.7% accuracy. Adding the FSC significantly boosts performance to 77.0%, indicating the advantage of incorporating frequency-domain information. Further introducing the MWTF module improves accuracy to 79.3%, demonstrating its ability to capture fine-grained temporal-frequency dynamics. A slight improvement is also observed when applying 3D coordinate normalization across different configurations, suggesting it can further stabilize input distributions and facilitate downstream spiking representation learning. Overall, these results confirm the progressive effectiveness of the proposed modules, with MWTF contributing most significantly on top of FSC.
4.4.2. Ablation study on the FSC module
Table 3. Ablation study on the FSC module.
Configuration $u$ Interaction Window Acc (%)
Without FSC 70.5
Real branch only - Learnable 72.3+1.8
Imaginary branch only - Learnable 72.1-0.2
Real + Imag (no fusion) - Learnable 73.6+1.5
Only $u=1$ branch $u=1$ Learnable 75.4+1.8
Only $u=2$ branch $u=2$ Learnable 74.9-0.5
Full FSC $u=1,2$ Learnable 78.3+3.4
Full FSC $u=1,2$ Hamming 78.0-0.3
Full FSC $u=1,2$ None 77.8-0.2
To validate the effectiveness of the proposed FSC module in enhancing spiking feature representation, we conduct comprehensive ablation studies, as shown in Table 3. Introducing only the real or imaginary branch improves performance from 70.5% to 72.3% and 72.1%, respectively, suggesting that both components contribute meaningfully. Combining them without interaction ($u$) further increases accuracy to 73.6% (+1.5%), confirming their complementarity. Enabling single interaction branches—either $u=1$ or $u=2$—leads to 75.4% and 74.9%, respectively, while using both ($u=\{1,2\}$) achieves the best performance at 78.3%, yielding a +3.4% gain over the best single-branch configuration and +7.8% over the baseline. Replacing the learnable window with a fixed Hamming window slightly reduces accuracy to 74.9%, and removing the window entirely yields 77.8%, showing the window’s limited but non-negligible contribution. These results demonstrate that the core benefit of FSC arises from complex-valued modeling and frequency-domain interaction mechanisms. To further understand the role of the learnable window, we conduct a visualization analysis across FSC layers, as detailed in Appendix D.2.
4.4.3. Qualitative Visualization of FSC module
To better understand the spiking neural behavior of our model, we visualize the joint-time ($V\times T$) spiking heatmaps from a representative “drinking water” sample, as shown in Figure 5. For each of the four FSC layers, we display activations from the real, imaginary, and fused branches, where grayscale intensity encodes the summed spike responses over the channel dimension.
Figure 5. Spike response visualizations of FSC for the “drinking water” action sample.
Figure 6. Visualization of upsampled detail coefficients $\hat{D}_{j}$ and scaling coefficients $\hat{S}_{j}$ across levels $j=0,1,2$, along with the final approximation coefficient $S_{3}$ in MWTF. Each heatmap shows the spike response intensity over joints and time steps, revealing the multi-scale temporal characteristics captured by the spiking cross-attention mechanism.
The fused features consistently exhibit smoother and more structured spatiotemporal patterns, indicating that combining real and imaginary branches enhances feature stability and expressiveness. In contrast, the real branch shows sparser, localized firing, while the imaginary branch captures more dispersed activity. Importantly, in the final FSC layer, joints associated with the right arm—specifically the right elbow (10), wrist (11), and hand (12)—exhibit prominent activations, aligning with the expected motion of lifting a cup. Symmetric activity is also observed in joints 6–8 (left arm), reflecting bilateral arm coordination. These observations validate that FSC learns semantically meaningful joint dynamics, and that the fused representation captures discriminative patterns critical for accurate action recognition.
4.4.4. Ablation study on the MWTF module.
To further analyze the effect of the MWTF module, we conduct an ablation study on its two key hyperparameters: $k$ (number of frequency coefficients) and $J$ (number of decomposition levels). The results, presented in Table 4, evaluate the trade-off between model complexity and classification accuracy.
@@ 4.4.3. Qualitative Visualization of FSC module
Table 4. Comparison of experimental results under different MWTF settings ($k$ and $J$)
MWTF $k$ $J$ Param. (M) FLOPs/SOPs (G) Acc (%)
- - - 1.65 1.60/0.312 78.3
4 1 1.68 1.61/0.313 78.9+0.6
8 1 1.73 1.62/0.314 79.8+0.9
8 2 1.74 1.62/0.315 80.3+0.5
8 3 1.74 1.62/0.317 80.5+0.2
16 3 1.83 1.64/0.320 80.5-
Starting from the baseline model without MWTF ($78.3\%$), incorporating low-order configurations ($k{=}4,J{=}1$) leads to a modest gain ($+0.6\%$), with the best performance ($80.5\%$) achieved at $k{=}8,J{=}3$. This validates the importance of multi-level frequency decomposition. However, further increasing $k$ to 16 brings no improvement, indicating saturation.
Table 5. Ablation study on spiking cross-attention and inclusion of $S_{J}$ in MWTF ($k{=}8,J{=}3$).
Setting $\hat{D}_{j}$ Branch $\hat{S}_{j}$ Branch $\hat{S}_{J}$ Acc. (%)
None (Baseline) 78.3
Only $\hat{D}_{j}$ 78.9+0.6
Only $\hat{S}_{j}$ 78.7-0.2
Only $S_{J}$ 78.4-0.3
$\hat{D}_{j}$ + $\hat{S}_{j}$ 79.8+0.4
$\hat{S}_{j}$ + $S_{J}$ 79.2-0.6
$\hat{D}_{j}$ + $S_{J}$ 79.4+0.2
Full (Ours) 80.5+1.3
To further dissect MWTF’s contributions, we ablate the two spiking cross-attention branches ($\hat{D}_{j}$, $\hat{S}_{j}$) and the use of the final approximation coefficient $\hat{S}_{J}$. Results in Table 5 show that each component brings incremental benefits, with $\hat{D}_{j}$ Branch alone improving performance to $78.9\%$, and the full configuration reaching $80.5\%$. To complement the analysis, we also investigate the evolution of the learnable fusion weight $\beta$ training. The results, including its stabilization behavior, are provided in Appendix D.1.
4.4.5. Interpretability Analysis of MWTF via CAM Visualization
To further validate the interpretability of the MWTF module, we visualize the joint-level Class Activation Maps (CAMs) derived from its output features for three representative actions: Drink Water, Stand Up,Reach into Pocket,Phone Call,Tear up Paper and Throw(Figure 7). These CAMs highlight the spatial regions most responsive in the model’s representation. In the Drink Water example, high activation appears near the right wrist and elbow, matching the expected movement pattern. In Stand Up, the model emphasizes lower body joints, while Reach into Pocket shows localized attention around the hip and wrist regions. These joint-specific activations are consistent with the underlying motion characteristics of each action. The results demonstrate that MWTF’s multi-scale frequency decomposition effectively captures action-relevant joint responses, confirming the strength of time-frequency analysis in spiking skeleton-based action recognition.
4.4.6. Qualitative Visualization of Multi-Scale Coefficients
To assess the impact of multilevel decomposition and spiking cross-attention in MWTF, we visualize the upsampled detail $\hat{D}_{j}$ and scaling $\hat{S}_{j}$ coefficients ($j=0,1,2$), along with the final approximation coefficient $S_{3}$, for a representative drinking water action (Figure 6). The heatmaps reveal that $\hat{D}_{j}$ emphasizes localized, high-frequency joint activations, while $\hat{S}_{j}$ captures smoother, global trends. Particularly, $S_{3}$ exhibits temporally stable responses, reflecting semantic patterns preserved through hierarchical decomposition. The complementarity between $\hat{D}_{2}$ and $\hat{S}_{2}$ further demonstrates the effectiveness of spiking cross-attention in disentangling multi-scale temporal dynamics for robust recognition. A more comprehensive spectral analysis based on joint-wise frequency decomposition is provided in Appendix D.3.
(a) Drink Water
(b) Stand Up
(c) Reach into Pocket
(d) Phone Call
(e) Tear Up Paper
(f) Throw
Figure 7. Class Activation Maps (CAMs) based on MWTF outputs for three representative actions. Warmer colors indicate higher joint-level importance.
5. Conclusion